• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name jpchen

    Log entry time 16:46:52 on July26,2006

    Entry number 175394

    keyword=Cryotarget all set for beam up to 10 uA

    Dave Meekins completed swtich of Lh2 loops from loop1 (which has Pb foil in) to loop 2 at around 1:30 pm.
    The temperature now for LH2 is 19K and LD2 is 22K.
    Pressure is about 30 psi for both LH2 and LD2 loops (bottle pressure reading).
    LH2 fan is 60Hz and LD2 is 48 Hz. With about 80W HP heater in each loop, the coolant flow is about 10 g/s, while Hall C is 12.5 g/s.
    We should be all set for beam and A+C is comfortably below 25 g/s.